An electron and proton enter a magnetic field perpendicularly. Both have same kinetic energy. Which of the following is true? ·

(A ) Trajectory of electron is less curved (B ) Trajectory of proton is less curved (C ) Both trajectories a1·e equally curved (D ) Both move on straight line path 

the answer given is A but i think it should be B

Grade:Select Grade

1 Answers

Gopal Gopakumar
7 years ago
Here, less curved just means that radius is very high.
 
Given that Kinetic energy is same..
=> 1/2*mv= K (say)
with some rearrangement,
v = (2K/m)1/2  ….
=> m*v = (2km)½.…
Now, radius r = mv/qb
 
here magnitude of q is same for electron and proton and magnetic field is same.
which means radius is proportional to mv .
=> radius is proportional to (2Km)1/2
since proton has more m, it will have a larger radius and hence its path will be less curved.